Divide the number 50 into two parts so that if one part is divided by 6 and the other part is divided by 8, the sum of the answers will be 9.
Answer by checkley71(8403)

x & (50-x)
x/6+(50-x)/8=9
(8x+6[50-x])/6*8=9
(8x+300-6x)/48=9 cross multiply
2x+300=432
2x=432-300
2x=132
x=132/2
x=66 answer.
50-66=-16 answer.
66/6+(50-66)/8=9
11-16/8=9
11-2=9
9=9